OpenStudy (jskinny0831):

Suppose you drop a ball from a window 32 meters above the ground. The ball bounces to 80% of its previous height with each bounce. What is the total number of meters (rounded to the nearest tenth) the ball travels between the time it is dropped and the 10th bounce?

OpenStudy (jskinny0831):

@surjithayer

OpenStudy (jskinny0831):

i got 253.64 by doing \[s _{10}= 2* 32(1-0.80^10)/1-0.80 -32 \]

OpenStudy (jskinny0831):

A. 110.8 meters B. 277.1 meters C. 253.6 meters D. 221.6 meters

OpenStudy (anonymous):

a=32,r=80/100=4/5 10 th bounce means 20 terms \[s=32*\frac{1- .8^{20} }{ 1-0.8 }\]

OpenStudy (jskinny0831):

is 253.6 correct?

OpenStudy (jskinny0831):

@surjithayer

OpenStudy (jskinny0831):

@mathmale i just wanna know if i did it right, please let me know

OpenStudy (jskinny0831):

@soprano.h.d0816

OpenStudy (anonymous):

 i am not clear if it is 19 exponential power or 20 exponential power.

OpenStudy (jskinny0831):

where did you get 19 and 20 from? its asking for the 10th bounce

OpenStudy (anonymous):

one coming down and one going up.

OpenStudy (jskinny0831):

ohhhh okay i get it now thank you

OpenStudy (anonymous):

i am leaving now.